OpenStudy (anonymous):

(radical 7 + radical 2) x (radical 7 - radical 2)

OpenStudy (anonymous):

7-2=5

OpenStudy (anonymous):

(a+b)(a-b)=a^2-b^2 use this rule... here a=sqrt(7) b=sqrt(2)